### v1

# Hyp-OW: Exploiting Hierarchical Structure Learning with Hyperbolic Distance Enhances Open World Object Detection

Abstract
Open World Object Detection (OWOD) is a challenging and realistic task that extends beyond the scope of standard Object Detection task. It involves detecting both known and unknown objects while integrating learned knowledge for future tasks. However, the level of ’unknownness’ varies significantly depending on the context. For example, a tree is typically considered part of the background in a self-driving scene, but it may be significant in a household context. We argue that this external or contextual information should already be embedded within the known classes. In other words, there should be a semantic or latent structure relationship between the known and unknown items to be discovered. Motivated by this observation, we propose Hyp-OW, a method that learns and models hierarchical representation of known items through a SuperClass Regularizer. Leveraging this learned representation allows us to effectively detect unknown objects using a Similarity Distance-based Relabeling module. Extensive experiments on benchmark datasets demonstrate the effectiveness of Hyp-OW achieving improvement in both known and unknown detection (up to 6 points). These findings are particularly pronounced in our newly designed benchmark, where a strong hierarchical structure exists between known and unknown objects.
1 Introduction
Advances in Object Detection (OD) have unlocked a plethora of practical applications such as robotics  Zhou et al. (2022), self-driving cars  Balasubramaniam and Pasricha (2022), manufacturing  Malburg et al. (2021), and medical analysis  Yang and Yu (2021). Recent breakthroughs in attention-based neural network architecture, such as Deformable Transformers  Zhu et al. (2021), have yielded impressive performance in these settings. However, most of these approaches assume a fixed number of classes (closed-world assumption), which is rare in reality. Continual Object Detection  Menezes et al. (2023) takes a step further by incrementally adding new classes, resulting in a distribution shift in the input and the well-known phenomenon of catastrophic forgetting(Kirkpatrick et al., 2017; Doan et al., 2021) where the network forgets previously learned knowledge. Open World (OW) Bendale and Boult (2015) takes these assumptions even further, introducing the detection and integration of newly discovered classes.
While the seminal work by  Bendale and Boult (2015) introduced the Open World (OW) framework, further advancements by  Joseph et al. (2021) extended it in two key aspects: the detection task and continual learning. However, a significant challenge within this framework lies in the absence of annotations for unknown objects, leading to biases towards known labels and potential confusion between unknown items and the background. This bias significantly impedes the accurate identification of unknown objects and presents a major hurdle in the detection process.
We can summarize previous attempts to solve this problem into three main categories. The first category includes works such as Joseph et al. (2021); Gupta et al. (2022); Zohar et al. (2022), which relied on a learned objectness score to relabel the background as potential unknowns. Another direction of research focused on clustering classes to better isolate unknowns (Wu et al., 2022b; Yu et al., 2022). Additionally, some works, like Kim et al. (2022); Wu et al. (2022a), introduced a decoupled approach where the classification and localization heads are separated. The objective is to remove label information and instead capture the shared features that makes these labelled items relevant as objects.
However, these works fail to address a crucial problem, which is defining what constitutes an "unknown" object. Currently, there is no clear definition or prior knowledge available to effectively distinguish unknowns from the background. Its interpretation greatly varies depending on the context. For example, in a driving scene, a "debris on the road" could be considered an unknown object  Balasubramaniam and Pasricha (2022), while in a camera surveillance context, it might be perceived as part of the background  Ingle and Kim (2022). Without considering the context, these works can only learn to differentiate knowns and unknowns at low level features such as texture or shape. As a consequence, they miss to model any hierarchical structures and similarities between known and unknown items, whether at the image level or dataset level.
Acknowledging this context information, we argue that a hierarchical structural relationship must exist between the objects to be discovered and the known items  Hosoya et al. (2022). This hierarchy is characterized by classes that share the same semantic context, belonging to the same category such as vehicles, animals, or electronics. Such hierarchical relationships enable the retrieval of common features and facilitate the discovery of unknown objects. For instance, a model trained on objects related to driving scenes can adequately detect stop signs or traffic lights but is not expected to recognize unrelated objects like a couch or any furniture.
In light of this discrepancy, we propose to learn hierarchical relationships between items in order to effectively utilize the representation of known objects for the discovery of unknown items. Ideally, items belonging to the same family (or category) should be closer to each other while being further away from different families (e.g., animals versus vehicles). To capture these structures, Hyperbolic Distance (Nickel and Kiela, 2018; Park et al., 2021), which naturally maps hierarchical latent structures, such as graphs or trees, emerges as a natural distance metric. This mapping in the hyperbolic space exhibits the desirable property of capturing the affinity between unknown items and known items, thereby enhancing the detection of unknown objects.
Figure 1: t-SNE plot of the learned class representations, with colors representing their respective categories. Our SuperClass Regularizer (right) enhances the hierarchical structure by grouping together classes from the same category while pushing apart classes from different categories.
Contribution
Motivated by the aforementioned literature gap, we propose a Hyperbolic Distance-based Adaptive Relabeling Scheme for Open World Object Detection (dubbed Hyp-OW). Our contribution can be summarized in three parts:
  • Hyp-OW is a simple yet effective method that learns inherent hierarchical structure between objects grouping item from the same category closer while pushing classes from different categories further apart through a SuperClass Regularizer (illustrated in Figure 1, right).
  • We propose an Adaptive Relabeling Scheme that enhances the detection of unknown objects by leveraging the semantic similarity between known and unknown objects in the hyperbolic space.
  • Our experiments demonstrate significant improvements in both unknown detection (up to $6\%$) and known accuracy performance (up to $5\%$) with Hyp-OW . These gains are particularly prominent when evaluating on our (designed) Hierarchical dataset, highlighting the advantages of our method in the presence of high inherent hierarchical structures.
2 Related Work
2.1 Open World Object Detection
The OWOD framework, introduced by Joseph et al. (2021), has inspired many recent works due to its realistic and close-to-real-world setting that integrates newly discovered items into the base knowledge progressively. While the first stream of work was originally based on the Faster-RCNN model (Joseph et al., 2021; Yu et al., 2022; Wu et al., 2022b, a), more recent works have utilized Deformable Transformers due to their superior performance (Gupta et al., 2022; Zohar et al., 2022). Joseph et al. (2021) introduced ORE, a Faster-RCNN-based model that learns class prototypes using contrastive learning with Euclidean distance. However, their approach relied on a held-out validation set where unknown items are explicitly labeled to learn an energy-based model to discriminate unknown items. Yu et al. (2022) extended this setting by minimizing the overlap between the distributions of unknown and known classes. OW-DETR (Gupta et al., 2022) designed a novelty-branch head to relabel the top-k highest background scores as unknowns. These pseudo-labels relied on unmatched bounding box proposals with high backbone activation being selected as unknown objects. On the other hand, Wu et al. (2022a) decoupled the localization and classification tasks (introduced by (Kim et al., 2022)) by learning a class-free head to localize objects. Recently, PROB (Zohar et al., 2022) learned a probabilistic objectness score by learning common statistics for all objects using Mahalanobis distance (Lee et al., 2018) and considered all the remaining bounding box proposals as unknown items. During the evaluation phase, they filter out proposal bounding boxes using the latter probabilistic models.
2.2 Class-Agnostic Object Detection
Another stream of work in the field of object detection is dubbed class-agnostic object detection, which focuses on localizing objects  (Kim et al., 2022; Wu et al., 2022a; Jaiswal et al., 2021). The objective is to remove the class label information and learn a shared low-level feature representation that effectively captures the essence of an object.  Kim et al. (2022) designed a pure localization head by introducing a second branch that is decoupled from the classification head.  Jaiswal et al. (2021) introduced an adversarial objective loss function that penalizes label information in the encoded features. Pixel-wise class-free object detection  Gonçalves et al. (2022) used texture gray level quantization to retrieve objects. Saito et al. Saito et al. (2022) designed a new data augmentation method that pastes an annotated object onto an object-free background.  Maaz et al. (2022) leveraged language models to improve unknown detection with their Multi-Modal Vision Transformers.
2.3 Learning hierarchical Representation with Hyperbolic Distance
Pointcare embeddings have been widely used in the literature to learn hierarchical structures from complex symbolic or multi-relational data, which can be represented by graphs or trees, such as social networks or taxonomies  (Nickel and Kiela, 2018; Law et al., 2019). Due to its good performance, it has been applied to image classification as well  (Khrulkov et al., 2020a; Yan et al., 2021; Yue et al., 2023; Ermolov et al., 2022). For example,  Yan et al. (2021) used hierarchical clustering to approximate a multi-layered tree structure representation that guides the hyperbolic distance learning process. Similarly, Liu et al. (2020) used taxonomy embedding from GloVe  (Pennington et al., 2014) to learn a finer-grained representation. Hyperbolic distance has also been used for object detection  (Lang et al., 2022; Ge et al., 2022).  Ge et al. (2022) was interested in learning context-object association rules by reasoning on different image scales. However, none of them leveraged the learned hyperbolic distance to retrieve unknowns items for OWOD.
3 Background
3.1 Problem Formulation
OWOD framework describes the setting where a user receives over the time a stream of $T$ tasks indexed by $t\in[1,T]$. Every task $t$ contains $C_{t}\in\mathbb{N}^{*}$ known classes ( denoted by set $\mathcal{K}^{t}$111whenever there is no ambiguity we will remove the task index $t$ to de-clutter the notation). The goal is to train an object detector module $f$ to accurately recognize the known classes but also discovering unknown classes (denoted by set $\mathcal{U}^{t}$). At the end of task $t$, $\Delta_{t}$ unknown classes are labelled (with an oracle) and included in the next task $t+1$ (containing now $\sum_{j=1}^{t}\Delta_{j}=C_{t}$ known classes). The process repeat until task $T$ that does not contain anymore unknowns.
The dataset of task $t$ is defined as $\mathcal{D}^{t}=\{\mathcal{I}^{t},\mathcal{Y}^{t}\}$ where $\mathcal{I}^{t}=\{\mathbf{I}_{t,1},\mathbf{I}_{t,2},...,\mathbf{I}_{t,N_{t}}\}$, are image inputs and $\mathcal{Y}^{t}=\{\mathbf{y}_{t,1},\mathbf{y}_{t,2},...,\mathbf{y}_{t,N_{t}}\}$ are the corresponding labels (There are $N_{t}$ images for task $t$). Each image $I_{t,i}$ contains a set of annotations $\mathbf{Y}_{t,i}=[\mathbf{l}_{t,i},\mathbf{x}_{t,i},\mathbf{y}_{t,i},\mathbf{w}_{t,i},\mathbf{h}_{t,i}]$ where $\mathbf{l}_{t,i}\in\{0,1\}^{C}$ denotes the object classes and $[\mathbf{x}_{t,i},\mathbf{y}_{t,i},\mathbf{w}_{t,i},\mathbf{h}_{t,i}]$ are the bounding box coordinates. Throughout the training, we will be storing item in a replay buffer $\mathcal{M}$ with a capacity of $m$ exemplar per class. We denote $\mathcal{B}$ the incoming batch.
We follow the setting of OWOD Joseph et al. (2021) where a set of $K$ examplars of each class is stored in a replay buffer at the end of each task $t$ (to mitigate forgetting). Those examplars are then replayed, i.e after task $2$, we replay $(2\cdot K\cdot 20)$ examplars both coming from task $1$ and $2$ ($K$ examplars from each task and $20$ classes per task).
3.2 Deformable Transformers for OWOD
We adopt Deformable Transformers (Zhu et al., 2021) as our base detector, as suggested by Gupta et al. (2022), due to its simplicity and high performance. This architecture processes the image input through a set of encoder-decoder modules to produce $Q$ queries output embeddings $\{\mathbf{q_{i}}\}_{i=1}^{Q}$, where $\mathbf{q_{i}}\in\mathbb{R}^{d}$ ($d$ being the query embedding dimension). These queries are then passed to the bounding box and classification heads, which respectively localize the labeled items and predict their classes. A bipartite matching algorithm (specifically, the Hungarian algorithm (Kuhn, 1955)) is used to match the labeled ground-truth items with each query. The remaining challenge is then to determine whether unmatched queries contain potential unknowns.
3.3 Hyperbolic Embeddings
A Hyperbolic space is a $n$-dimensional Riemann manifold defined as $(\mathbb{B}_{c}^{n},g^{\mathbb{M}})$ with its Poincare ball $\mathbb{B}_{c}^{n}=\{x\in\mathbb{R}^{n}:c\lVert x\rVert^{2}\leq 1,c\geq 0\}$ ($c$ being the constant curvature) and equipped with a Riemannian metric $g^{\mathbb{M}}=\lambda_{x}^{2}g^{E}$ where $\lambda_{x}^{c}=\frac{2}{1-c\lVert x\rVert^{2}}$, $g^{E}=\mathbf{I}_{n}$ is the Euclidian metric tensor. The transformation from the Euclidian to hyperbolic space is done via a bijection termed exponential mapping $\exp_{\mathbf{b}}^{c}:\mathbb{R}^{n}\rightarrow\mathbb{B}_{c}^{n}$.
$\exp_{\mathbf{b}}^{c}(\mathbf{x})=\mathbf{b}\otimes_{c}(\tanh{(\sqrt{c}\frac{\lambda_{\mathbf{b}}^{c}\lVert\mathbf{x}\rVert}{2})}\frac{\lVert\mathbf{x}\rVert}{\sqrt{c}\lVert\mathbf{x}\rVert})$ (1)
where $\mathbf{b}$ represents the base point. The latter is often empirically taken as $\mathbf{b}=\mathbf{0}$ to simplify the formulas without impacting much the results (Ermolov et al., 2022). We will also adopt this value in our study.
Inside this hyperbolic space, the distance between two points $\mathbf{x},\mathbf{y}\in\mathbb{B}_{c}^{n}$ is computed as:
$d_{hyp}(\mathbf{x},\mathbf{y})=\frac{2}{\sqrt{c}}\arctan{(\sqrt{c}\lVert-\mathbf{x}\otimes_{c}\mathbf{y}\rVert)}$ (2)
where the addition operation $\otimes_{c}$ is defined as : $\mathbf{x}\otimes_{c}\mathbf{y}=\frac{(1+2c\left<\mathbf{x},\mathbf{y}\right>+c\lVert\mathbf{y}\rVert^{2})\mathbf{x}+(1-c\lVert\mathbf{x}\rVert^{2})\mathbf{y}}{1+2c\left<\mathbf{x},\mathbf{y}\right>+c^{2}\lVert\mathbf{x}\rVert^{2}\lVert\mathbf{y}\rVert^{2}}$.
From now on, we will denote $\mathbf{z_{i}}$ the projection of the queries $\mathbf{q_{i}}$ into the hyperbolic embedding space, i.e, $\mathbf{z_{i}}=\exp_{\mathbf{b}}^{c}(\mathbf{q_{i}})$. Note that when $c=0$, the distance boils down to the cosine similarity defined as:
$D_{cos}(\mathbf{x},\mathbf{y})=2-2\frac{<\mathbf{x},\mathbf{y}>}{\lVert\mathbf{x}\rVert_{2}\cdot\lVert\mathbf{y}\rVert_{2}}$ (in this case no exponential mapping is needed, i.e $\mathbf{z_{i}}=\mathbf{q_{i}}$222from now on, we will refer to $\mathbf{z_{i}}$ interchangeably for both the query and its corresponding embedding $\mathbf{q_{i}}$)
4 Hyp-OW
In this section, we provide a detailed explanation of each module of our proposed method. Hyp-OW can be summarized by three main components (Figure 2): a Hyperbolic Metric Distance learning (Section 4.1), a SuperClass Regularizer (Section 4.2) and an Adaptive Relabeling Scheme to detect unknowns (Section 4.3).
Figure 2: Overview of each component of Hyp-OW.There are three main components: the Hyperbolic Contrastive Loss, which learns a hierarchical structural representation of each class; the SuperClass Regularizer, which models the semantic relationship among classes to ensure proximity within the same category and distance from different categories; and the Adaptive Relabeling module, which utilizes the hierarchical structure to compute the hyperbolic distance between candidate proposals and known items. If this distance $d$ is lower than a certain threshold ($\delta$), the proposal is relabelled as unknown.
4.1 Metric Learning with Hyperbolic Distance
We learn feature representation in the hyperbolic embedding space using a contrastive loss. The idea is to move closer features belonging to the same class $c$ while repetlling then away from any features from class $j\neq c$. Let’s denote $\mathbf{z_{i}^{c}}$ any query $i$ matched with class $c\in\mathcal{K}$333by abuse of notation, we will also use $c$ for the class label since the meaning can be infered from the context. During the training, we maintain a replay buffer $\mathcal{M}$ where we store $m$ embedding features per class.
For every query element $\mathbf{z_{i}^{c}}$ of the incoming batch $\mathcal{B}$, we sample $k=1$ elements of the same class $c\in\mathcal{K}$ from the replay buffer $\mathcal{M}$ (which will be considered as the positive examples, the $22|\mathcal{B}|-$ 2 remaining samples be considered as the negative examples).
For simplicity we denote all the elements from the buffer and the batch $\mathcal{A}=\mathcal{B}\cup\mathcal{M}$. Every element $\mathbf{z_{i}^{c}}\in\mathcal{B}$ (respectively $i\in\mathcal{M}$) has its positive counterpart examples $\mathbf{z_{i^{+}}}$ of same class $c$ from $\mathcal{M}$ (respectively from $\mathcal{B}$) and $|\mathcal{A}|-2$ negative examples (both from $\mathcal{B}$ and $\mathcal{M}$) denoted as $\mathbf{z_{i^{-}}}$.
Defining a temperature $\tau_{1}$, the contrastive loss is then expressed as:
$\mathscr{L}_{hyp}=-\displaystyle{\sum_{i\in\mathcal{A},c\in\mathcal{K}}}\log\frac{\exp(\frac{-d_{hyp}(\mathbf{z_{i}^{c}},\mathbf{z_{i^{+}}})}{\tau_{1}})}{\displaystyle{\sum_{i^{-}\in\mathcal{A}\setminus\{i,i^{+}\}}}\exp(\frac{-d_{hyp}(\mathbf{z_{i}^{c}},\mathbf{z_{i^{-}}})}{\tau_{1}})}$ (3)
This loss aims at attracting representation of $\mathbf{z_{i}^{c}}$ closer to its positive counterpart $\mathbf{z_{i^{+}}}$ while repelling from other classes representation $\mathbf{z_{i^{-}}}$,$i\in\mathcal{A},c\in\mathcal{K}$.
4.2 SuperClass Regularization
Many real-world datasets exhibit a natural hierarchical structure, where classes can be organized into categories. For instance, dogs and cats belong to the animal category, while cars and trucks belong to vehicles category. To leverage this hierarchical information, we propose a SuperClass Regularizer (we will use SuperClass and category interchangeably throughout this study), which encourages classes within the same category to be closer in the embedding space while pushing them away from classes in different categories.
Let’s denote $\mathcal{S}_{p}$ the set of class indexes belonging to Category $p=1...P$. We approximate the category $p$ embedding by computing the Hyperbolic Average (Khrulkov et al., 2020b) (dubbed HypAve) of every embedding $\{\mathbf{z_{i}^{c}}\}_{i\in\mathcal{M}}$ of classes $c\in\mathcal{S}_{p}$ from the buffer $\mathcal{M}$, that is:
$HypAve(\{\mathbf{z_{i}^{c}}\}_{i\in\mathcal{M},c\in\mathcal{S}_{p}})=\frac{\displaystyle{\sum_{i\in\mathcal{M},c\in\mathcal{S}_{p}}}\gamma_{i}\mathbf{z_{i}^{c}}}{\displaystyle{\sum_{i\in\mathcal{M}}}\gamma_{i}}$ (4)
where $\gamma_{i}=\frac{1}{\sqrt{1-c\lVert x_{i}\rVert^{2}}}$ are the Lorentz factors. To de-clutter the notation, we will denote $\mathbf{\overline{z}_{p}}=HypAve(\{\mathbf{z_{i}^{c}}\}_{i\in\mathcal{M},c\in\mathcal{S}_{p}})$ the Hyperbolic Average feature representation of category $p$. For every element $\mathbf{z_{i}^{c}}$ of a batch $\mathcal{B}$, we sample its category embedding $\mathbf{\overline{z}_{p}}(c\in\mathcal{S}_{p})$ from the buffer $\mathcal{M}$ and use the same contrastive loss:
$\mathscr{L}_{reg}=\displaystyle{\sum_{i\in\mathcal{A},c\in\mathcal{S}_{p}}}-\log\frac{\exp(\frac{-d_{hyp}(\mathbf{z_{i}^{c}},\mathbf{\overline{z}_{p}})}{\tau_{2}})}{\displaystyle{\sum_{k\neq p}}\exp(\frac{-d_{hyp}(\mathbf{z_{i}^{c}},\mathbf{\overline{z}_{k}})}{\tau_{2}})}$ (5)
This loss encourages the features $\mathbf{z}_{i}^{c}$ of each class $c$ to be closer to its corresponding category embedding $\mathbf{\overline{z}}_{p}$, while simultaneously pushing it away from embeddings of other categories $\mathbf{\overline{z}}_{k}$, $k\neq p$.
4.3 Adaptive Relabeling of Unknowns with Hyperbolic Distance
We introduce our Adaptive Relabeling module, which dynamically adapts to the batch statistics to effectively retrieve unknown items (leveraging the learned hyperbolic embedding). Recall for the classification task, we have two types of queries: matched queries with ground-truth labels (found with the Hungarian Algorithm), and unmatched queries that may contain unknown items. The matched queries will establish a threshold criterion to select unknown queries from the unmatched set.
[…]
$HypAve(\{\mathbf{z_{i}^{c}}\}_{i\in\mathcal{M}})=\frac{\displaystyle{\sum_{i\in\mathcal{M}}}\gamma_{i}\mathbf{z_{i}^{c}}}{\displaystyle{\sum_{i\in\mathcal{M}}}\gamma_{i}}$ which can be seen as the centroid of each class $c$ in the hyperbolic embedding space.
[…]
Overall loss
All the aforementioned losses are finally optimized together as:
$\mathcal{L}=\mathcal{L}_{cls}+\mathcal{L}_{bbox}+\alpha\mathcal{L}_{hyp}+\beta\mathcal{L}_{reg}$ (7)
Where $\alpha,\beta\geq 0$ are coefficient controlling respectively the Hyperbolic and Regularizer importance.
5 Experiments
In this section, we begin by describing our experimental setup. We then present comparative results against benchmark baselines, followed by in-depth ablation analysis of each component of Hyp-OW . Due to space limitations, we will defer detailed information to the Appendix Section C and  D.
5.1 Experimental Setup
Datasets
We consider two benchmarks from the literature: the OWOD Split Joseph et al. (2021) and the OWDETR Split Gupta et al. (2022). While the latter (OWDETR Split) strictly separates superclasses across tasks the first (OWOD) has mild semantic overlap between knowns and unknowns across tasks (See Appendix Section A). To validate our hypothesis regarding the semantic relationship between knowns and unknowns, we introduce a third dataset called the Hierarchical Split. This dataset ensures that each task includes at least one class from each category, promoting a higher level of semantic similarity. This will be discussed in paragraph 5.2. Each dataset is defined by four tasks $t=1,2,3,4$, containing 20 labelled classes each, for a total of 80 classes. When task $t$ starts, only the label of classes belonging to that task are revealed. For instance, task $1$ only contains labels of classes from 0 to 19, while task $2$ only contains labels of classes from 21 to 39, and so on.
Implementation Details
We use Deformable DETR (Zhu et al., 2021) pretrained in a self-supervised manner (DINO Caron et al. (2021)) on Resnet-50 (He et al., 2016) as our backbone. The number of deformable transformer encoder and decoder layers are set to $6$. The number of queries is set to $Q=100$ with a dimension $d=256$. During inference time, the top-100 high scoring queries per image are used for evaluation. Additional detail are provided in Appendix Section B.
Metrics and Baselines
Following current metrics used for OWOD, we use mean average precision (mAP) for known items while U-Recall is the main metric used to quantify the unknown detection quality or each method Gupta et al. (2022); Wu et al. (2022a); Zohar et al. (2022); Maaz et al. (2022); Yu et al. (2022). Additional metric is discussed in Table 9 We consider the following baselines from literature: ORE-EBUI (Joseph et al., 2021), UC-OWOD (Wu et al., 2022b), OCPL (Yu et al., 2022), 2B-OCD (Wu et al., 2022a), OW-DETR (Gupta et al., 2022) and PROB (Zohar et al., 2022).
5.2 Benchmark Results
Dataset Similarity
To gain insights into the structure of each dataset, we introduce a semantic similarity measure based on GloVe’s embedding Pennington et al. (2014) (defined in Appendix Section A.1). This measure quantifies the similarity overlap between known and unknown items, with higher values indicating larger overlap. The three splits, OW-DETR Split (Low regime ), OWOD Split (Medium regime ), and Hierarchical Split (High regime ), exhibit a monotonic ranking in terms of similarity, with respective values of $0.27$, $0.33$, and $0.41$. This serves as a starting point for evaluating baseline methods under various scenarios.
Unknown Detection (U-Recall)
Table 1 shows the high performance gain of Hyp-OW over PROB on Medium regime and High regime of $3\%$ on average. This highlights the utility of learning hierarchical structured representations and retrieving unknowns based on their similarity with known objects, as opposed to PROB, which learns a single mean representation for all objects. For the Low regime our method is performing on-par with PROB except for task 1 which shows a surprising improvement of $6$ points. A possible explanation may come from the nature of the Object Detection (OD) task, where there can be significant overlap between bounding boxes. This encourages the model to learn classes that frequently co-occur, such as "person" and "backpack" or "teddy bear" and "bicycle" (See Figure 3). We provide qualitative and quantitative explanations in Appendix Section C.
Task 1 Task 2 Task 3 Task 4
U-Recall ($\uparrow$) mAP ($\uparrow$) U-Recall ($\uparrow$) mAP ($\uparrow$) U-Recall ($\uparrow$) mAP ($\uparrow$) mAP ($\uparrow$)
Low
ORE - EBUI 1.5 61.4 3.9 40.6 3.6 33.7 31.8
OW-DETR 5.7 71.5 6.2 43.8 6.9 38.5 33.1
PROB 17.6 73.4 (+0.7) 22.3 50.4 24.8 42.0 39.9
Hyp-OW (Ours) 23.9 (+6.3) 72.7 23.3 (+1.0) 50.6 25.4 46.2 (+4.2) 44.8 (+4.9)
Medium ORE - EBUI 4.9 56.0 2.9 39.4 3.9 29.7 25.3
UC-OWOD 2.4 50.7 3.4 8.7 16.3 24.6 23.2
OCPL 8.26 56.6 7.65 39.1 11.9 30.7 26.7
2B-OCD 12.1 56.4 9.4 38.5 11.6 29.2 25.8
OW-DETR 7.5 59.2 6.2 42.9 5.7 30.8 27.8
PROB 19.4 59.5 17.4 44.0 19.6 36.0 31.5
Hyp-OW (Ours) 23.5 (+4.1) 59.4 20.6 (+3.2) 44.4 26.3 (+6.7) 36.8 33.6 (+2.1)
High
OW-DETR 7.0 47.3 11.0 38.6 8.8 38.3 38.2
PROB 29.4 49.6 43.9 42.9 52.7 41.3 41.0
Hyp-OW (Ours) 34.9 (+5.5) 49.9 47.5 (+3.6) 45.5 (+2.6) 55.2 (+2.5) 44.3 (+3.1) 43.9 (+2.9)
Table 1: State-of-the-art comparison on the three splits for unknown detection (U-Recall) and known accuracy (mAP). Hyp-OW improves significantly the unknown detection (U-Recall) for the medium and high regime and known detection (mAP) for the low regime. Task 4 does not have U-Recall since all 80 classes are known at this stage.
Figure 3: Bounding boxes overlap. In the OD task, the high overlap between bounding boxes of frequently co-occurring objects can influence the model to learn their associations and correlations.
Known Accuracy (mAP)
From the known accuracy, Hyp-OW surpasses baseline benchmark on all tasks of Hierarchical Split and shows notable performance for the two last tasks of OW-DETR Split. This can be credited to the structural hierarchy learned that groups together class of same category (See t-SNE Figure 1 middle and right).
5.3 Ablation Analysis
We now propose an in-depth understanding of Hyp-OW by removing one by one each component and see its direct impact (illustrated quantitatively in Table 2 for Hierarchical Split).
Task 1 Task 2 Task 3 Task 4
U-Recall ($\uparrow$) mAP ($\uparrow$) U-Recall ($\uparrow$) mAP ($\uparrow$) U-Recall ($\uparrow$) mAP ($\uparrow$) mAP ($\uparrow$)
Hyp-OW (Ours) 34.9 49.9 47.5 45.5 55.2 44.3 43.9
w/ Cosine Distance 32.8 (-2.1) 49.0 (-0.9) 46.4 (-1.1) 45.4 55.4 43.2 (-1.1) 43.1
w/o SuperClass Regularizer 32.0 (-2.9) 50.0 47.1 45.1 52.9 (-2.3) 43.7 43.5
w/o Adaptive Relabeling 34.7 41.2 (-8.7) 47.6 38.9 (-6.6) 54.1 (-1.1) 36.5 (-7.8) 36.1 (-7.8)
Table 2: Impact of each component of Hyp-OW on hierarchical Split. We observe that the Relabeling module (third line) significantly reduces mAP while maintaining U-Recall. On the other hand, the SuperClass Regularizer and Cosine Distance have primarily an impact on unknown detection.
Adaptive Relabeling:
This module uses Eq 6 to relabel unmatched bounding boxes as unknowns. To assess the impact of this relabeling strategy, we compare it with an alternative approach used by PROB Zohar et al. (2022), where all unmatched queries are labeled as unknowns. Although the decrease in U-Recall is marginal, we observe a significant reduction in known accuracy (mAP). This performance degradation can be attributed to the over-prediction of patches as unknowns, which results in misclassification of known objects. Heatmaps in Figure 10 and  11 (Appendix) demonstrate the effectiveness of our module where we observe that unknowns belonging to the same category (sharing some color) as knowns exhibit lower Hyperbolic Distance.
Figure 4: Hyperbolic Category - Class Distance Heatmap. The SuperClass Regularizer (right) effectively separates different categories (left), as indicated by the increased distance between each animal class (bottom) and the vehicle, outdoor, and furniture categories (darker colors). Without this Regularizer (left), category inter-distance are much smaller (lighter color intensity).
SuperClass Regularizer:
By setting $\beta=0$, we no longer enforce the grouping of items from the same category in the hyperbolic space (compare t-SNE plots in Figure 1). As a result, we observe a reduction in U-Recall of $2.9$, $0.4$, and $2.3$ points, respectively (Table 2, second line). Heatmap Figure 4 illustrates the hyperbolic distance from each class to every category’s embedding (computed using Eq 4) with lighter colors indicating smaller distances. With our Regularizer (right plot), we observe a wider range of values spanning from $0.7$ to $2.30$, compared to a smaller range of $0.78$ to $1.2$ without the Regularizer. This highlights the effect of our Regularizer that pushes classes from different categories apart (shown as dark color in the right plot) while bringing classes of similar categories closer together. A more detailed plot can be found in Appendix Figure 12.
Cosine Distance:
Trading hyperbolic distance for the cosine distance ($c=0$) harms both the U-Recall and mAP as the hyperbolic embedding space is more suitable to learn data with latent hierarchical structure (highlighted by Figure 14 in Appendix).
6 Conclusion
The Open World Object Detection framework presents a challenging and promising setting, encompassing crucial aspects such as lifelong learning and unknown detection. In our work, we have emphasized the lack of a clear definition of unknowns and the need of a hierarchical or semantic relationship between known and unknown classes. This led us to propose Hyp-OW that focuses on learning and modeling the structural hierarchy within the dataset, which is then utilized for unknowns retrieval. Extensive experiments demonstrates significant improvement of Hyp-OW for both known and unknown detection (up to 6 points) particularly in the presence of inherent hierarchy between classes.
Acknowledgement
We would like to thank (in no particular order): Jiajing Guo, Shabnam Ghaffarzadegan, Yulian Guo, Sharath Gopal, Sanbao Bu and Jorge Ono for useful discussions and feedbacks throughout the work.
Appendix
The Appendix Section is organized as follow:
  • Section A provides information on the composition of each dataset and the introduction of our semantic similarity measure.
  • Section B presents the experimental details.
  • Section C includes comprehensive benchmark metrics results and visual plots (heatmaps) to illustrate our findings.
  • Section D discusses the results of the ablation analysis accompanied by visual figures for better understanding.
Appendix A Dataset Information
A.1 Dataset Composition
In this section, we provide the composition of each dataset split and explain the main differences between them. The splits are made with images taken from the PASCAL-VOC2007/2012 (Everingham et al., 2010) and MS-COCO dataset (Lin et al., 2014), containing a total of 80 classes grouped into 12 catogories (SuperClass): vehicle, outdoor, animal, accessory, sports, kitchen, food, furniture, appliance, electronic, indoor, and person. We categorize the different datasets into three regimes based on a semantic similarity metric (defined below), which quantifies the overlap between known and unknown items:
  • Low regime : OWDETR Split
  • Medium regime : OWOD Split
  • High regime : Hierarchical Split
The low regime implies no semantic similarity between classes of different tasks. For instance, OWDETR Split’s ( 3) classes of task 1 are composed of vehicle, animal, and person, which do not appear in tasks 2, 3, or 4. On the other hand, high regime shows high similarity between classes across all tasks. Each task of the Hierarchical split (Table 5) contains at least one class from each of the 12 categories. The OWOD Split is a trade-off between the aforementioned datasets.
Low regime : OWDETR Split
Task 1 Task 2 Task 3 Task 4
aeroplane traffic light frisbee laptop
bicycle fire hydrant skis mouse
bird stop sign snowboard remote
boat parking meter sports ball keyboard
bus bench kite cell phone
car chair baseball bat book
cat diningtable baseball glove clock
cow pottedplant skateboard vase
dog backpack surfboard scissors
horse umbrella tennis racket teddy bear
motorbike handbag banana hair drier
sheep tie apple tootbrush
train suitcase sandwhich wine glass
elephant microwave orange cup
bear oven broccoli fork
zebra toaster carrot knife
giraffe sink hot dog spoon
truck refrigerator pizza bowl
person bed donut tvmonitor
toilet cake bottle
sofa
Table 3: Composition of OWDETR Split. There is no overlap of categories between each task
Medium regime : OWOD Split
Task 1 Task 2 Task 3 Task 4
aeroplane truck frisbee bed
bicycle traffic light skis toilet
bird fire hydrant snowboard laptop
boat stop sign sports ball mouse
bottle parking meter kite remote
bus bench baseball bat keyboard
car elephant baseball glove cell phone
cat bear skateboard book
chair zebra surfboard clock
cow giraffe tennis racket vase
diningtable backpack banana scissors
dog umbrella apple teddy bear
horse handbag sandwhich hair drier
motorbike tie hot dog toothbrush
person suitcase broccoli wine glass
pottedplant microwave carrot cup
sheep oven hot dog fork
sofa toaster pizza knife
train sink donut spoon
tvmonitor refrigerator cake bowl
Table 4: Composition of OWOD Split. There is a mild overlap of categories between each task.
High regime : hierarchical Split
Task 1 Task 2 Task 3 Task 4
bicycle person bus truck
car motorbike train boat
traffic light aeroplane parking meter bench
fire hydrant stop sign cow zebra
bird horse elephant giraffe
cat sheep bear tie
dog umbrella handbag suitcase
backpack snowboard kite skateboard
frisbee sports ball baseball bat surfboard
skis cup baseball glove tennis racket
bottle sandwich fork spoon
wine glass orange knife bowl
banana broccoli carrot pizza
apple pottedplant hot dog donut
chair bed diningtable cake
sofa laptop remote toilet
tvmonitor mouse keyboard cell phone
microwave toaster sink refrigerator
oven clock scissors hair drier
book vase teddy bear toothbrush
Table 5: Composition of hierarchical Split. Each task contains at least one class of each category.
Quantifying Semantic Overlap between Knowns and Unknowns across Dataset
We propose a measure to quantify the semantic similarity overlap between known and unknown classes in each dataset. To achieve this, we utilize the GloVe (Pennington et al., 2014) embedding for each known class $c\in\mathcal{K}^{t}$ (and unknown class $k\in\mathcal{U}^{t}$), denoted by $\mathbf{\omega_{c}}$ ( respectively $\mathbf{\omega_{k}}$ ). For a given task $t\in[1,T]$, we define the semantic overlap between knowns and unknowns as:
$S_{t}=\frac{1}{|\mathcal{U}^{t}|}\displaystyle{\sum_{k\in\mathcal{U}^{t}}\max_{c\in\mathcal{K}^{t}}}\frac{<\mathbf{\omega^{c}},\mathbf{\omega^{k}>}}{\lVert\mathbf{\omega^{c}}\rVert_{2}\cdot\lVert\mathbf{\omega^{k}}\rVert_{2}},\forall t\leq T-1$ (8)
Higher value means higher similarity between the knowns and unknowns. For a given task $t$, this value calculates the similarity between each unknown and the highest-ranked known and averages it across all unknowns. Figure 5 shows the evolution of this similarity measure throughout the training. As expected, the three splits ranks monotonically as low similarity (OWDETR Split ), medium similarity (OWOD Split) and high similartity (Our hierarchical Split).
Figure 5: Semantic Similarity between knowns and unknowns across tasks for each Split.
Appendix B Implementation Details
In this section, we provide a recap of the definition and value of each hyperparameter used, as well as the learning rate schedule employed during training. The sequence of learning rate for each task is recalled in Table 7. Additional details are provided regarding the usage of the buffer in various loss functions.
Hyperparameters
Parameters Value Definition
$\alpha$ 0.05 Coefficient of the Hyperbolic Contrastive Loss Eq 3
$\beta$ 0.02 Coefficient of the SuperClass Regularizer Eq 5
$\tau_{1}$ 0.2 Temperature of the Hyperbolic Contrastive Loss
$\tau_{2}$ 0.4 Temperature of the SuperClass Regularizer
$m$ 10 Capacity of the replay buffer (examplar per class)
$c$ 0.1 curvature coefficient for the Hyperbolic Distance
$k$ 1 number of positive examples to be sampled for each anchor
batch size 3 N/A
GPUs 4 Nvidia RTX 3090 N/A
Table 6: Hyperparameters used for Hyp-OW .
hierarchical Split
Task $\mathcal{T}_{t}$ From Epoch to Learning rate $l_{r}$ Classes labeled
Task 1 0 to 40 $10^{-4}$ 0-19
Task 1 40 to 50 $10^{-5}$ 0-19
Task 2 50 to70 $10^{-4}$ 20-39
Task 2 fine-tuning 70 to 130 $10^{-4}$ 0-39
Task 3 130 to150 $10^{-4}$ 40-59
Task 3 fine-tuning 150 to210 $10^{-4}$ 0-59
Task 4 210 to 230 $10^{-4}$ 60-79
Task 4 fine-tuning 230-300 $10^{-4}$ 0-80 (no unknowns)
Table 7: Learning rate schedule throughout the 4 tasks.
Buffer Filling:
Throughout the training process, we store the embeddings of every encountered class $\mathbf{z_{i}^{c}}$ (where $c\in\mathcal{K},i\in\mathcal{B}$) which has a capacity of $m$ exemplars per class.
Buffer Interaction with Hyperbolic Contrastive Loss:
To compute the contrastive loss (Eq. 3), for each embedding $z_{i}^{c}$ in the batch, where $i\in\mathcal{B}$ and $c\in\mathcal{K}$ , we sample its positive counterpart $\mathbf{z_{i^{+}}}$ from the same class $c$ in the buffer.
Buffer Interaction with SuperClass Regularizer:
To compute the Hyperbolic Average $\mathbf{\overline{z}_{p}}$ ($p=1...P$) in the SuperClass Regularizer (Eq. 5), we sample from the buffer every embedding $\mathbf{z_{i}^{c}}$, $i\in M,c\in\mathcal{S}_{p}$ then use Eq 4 to calculate $\mathbf{\overline{z}_{p}}$.
Appendix C Additional Experimental Results
C.1 Detailed Results
Table 8 presents a detailed overview of the performance of all baselines across the three splits. It includes the mean Average Precision (mAP) of the Previous and Current known classes for tasks 2, 3, and 4. Hyp-OW exhibits a high level of plasticity, as evidenced by its accuracy on the current known classes (highlighted in the ’Current’ column of tasks 2, 3, and 4).This can be attributed to the hierarchical structure learned within the hyperbolic embedding space, which facilitates effective modeling and representation of the known classes.
Regime
Task 1 Task 2 Task 3 Task 4
U-Recall mAP ($\uparrow$) U-Recall ($\uparrow$) mAP ($\uparrow$) U-Recall ($\uparrow$) mAP ($\uparrow$) mAP ($\uparrow$)
Current Previous Current Both Previous Current Both Previous Current Both
Low
ORE - EBUI 1.5 61.4 3.9 56.5 26.1 40.6 3.6 38.7 23.7 33.7 33.6 26.3 31.8
OW-DETR 5.7 71.5 6.2 62.8 27.5 43.8 6.9 45.2 24.9 38.5 38.2 28.1 33.1
PROB 17.6 73.4 22.3 66.3 36.0 50.4 24.8 47.8 30.4 42.0 42.6 31.7 39.9
Hyp-OW (Ours) 23.9 (+6.3) 72.7 23.3 (+1.0) 59.8 42.2 50.6 25.4 49.3 39.8 46.2 (+4.2) 46.4 40.3 44.8 (+4.9)
Medium ORE - EBUI 4.9 56.0 2.9 52.7 26.0 39.4 3.9 38.2 12.7 29.7 29.6 12.4 25.3
UC-OWOD 2.4 50.7 3.4 33.1 30.5 31.8 8.7 28.8 16.3 24.6 25.6 15.9 23.2
OCPL 8.26 56.6 7.65 50.6 27.5 39.1 11.9 38.7 14.7 30.7 30.7 14.4 26.7
2B-OCD 12.1 56.4 9.4 51.6 25.3 38.5 11.6 37.2 13.2 29.2 30.0 13.3 25.8
OW-DETR 7.5 59.2 6.2 53.6 33.5 42.9 5.7 38.3 15.8 30.8 31.4 17.1 27.8
PROB 19.4 59.5 17.4 55.7 32.2 44.0 19.6 43.0 22.2 36.0 35.7 18.9 31.5
Hyp-OW (Ours) 23.5 $\boldsymbol{(+4.1)}$ 59.4 20.6 $\boldsymbol{(+3.1)}$ 51.4 37.4 44.4 26.3 $\boldsymbol{(+6.6)}$ 42.9 24.6 36.8 37.4 22.4 33.6 (+2.1)
High
OW-DETR 7.0 47.3 11.0 38.0 39.2 38.6 8.8 39.3 36.1 38.3 38.5 37.3 38.2
PROB 29.4 49.6 43.9 41.9 43.9 42.9 52.7 41.7 40.4 41.3 40.9 41.2 41.0
Hyp-OW (Ours) 34.9 (+5.2) 49.9 47.5 $\boldsymbol{(+3.5)}$ 42.0 49.0 45.5 $\boldsymbol{(+2.7)}$ 55.2 $\boldsymbol{(+2.5)}$ 44.4 44.1 44.3 $\boldsymbol{(+3.1)}$ 42.8 47.0 43.9 $\boldsymbol{(+3.0)}$
Table 8: Benchmark Results.
Next, we turn our attention to the issue of unknown object confusion and address it through the Absolute Open-Set Error (A-OSE) metric, as introduced by Joseph et al. (2021). The A-OSE quantifies the number of unknown objects that are incorrectly classified as known (lower is better). In Table 9, we present the A-OSE values alongside the U-Recall metric for the Hierarchical Split.
hierarchical Split Task 1 Task 2 Task 3
U-Recall ($\uparrow$) A-OSE($\downarrow$) U-Recall ($\uparrow$) A-OSE($\downarrow$) U-Recall ($\uparrow$) A-OSE($\downarrow$)
OW-DETR 7.0 42,540 11.0 26,527 8.8 20,034
PROB 29.4 14,962 43.9 8,929 52.7 5,387
Hyp-OW (Ours) 34.9 7,420 47.5 3,849 55.2 4,611
Table 9: Unknown Confusion. Hyp-OW achieves the highest U-Recall, indicating its superior ability to detect unknown objects. Additionally, it exhibits the lowest A-OSE, further emphasizing its capability to effectively differentiate between known and unknown objects.
C.2 Unveiling Co-occurrence Learning: The Influence of Bounding Box Overlap
In the Object Detection (OD) task, where bounding boxes can overlap, the model has a tendency to learn co-occurrence patterns of frequently appearing classes. This phenomenon is quantitatively demonstrated in the Heatmap Figure 6, where the model exhibits high similarity between classes that do not have any semantic relationship, such as ’person’ and ’tie’, or ’teddy bear’ and ’bicycle’ (highlighted in teal). This observation is further supported by Figures  7,  8 and  9, which provide qualitative examples of this co-occurrence learning.
Figure 6: Hyperbolic Knowns-Unknowns Distance Heatmap. Lighter colors indicate lower hyperbolic distance and higher similarity. In the presence of images with high bounding box overlap, the model can learn frequent associations, such as ’person’ and ’tie’ or ’backpack’ (teal hatched boxes, last row) or ’teddy bear’ and ’bicycle’.
Figure 7: Bounding boxes overlap between classes ’teddy bear’ and ’bicycle’.
Figure 8: Bounding boxes overlap between classes ’person’ and ’tie’.
Figure 9: Bounding boxes overlap between classes ’person’ and ’backpack’.
Appendix D Ablation Analysis
In this section, we present comprehensive results of the ablation analysis. We then delve into a detailed discussion of the impact of each component.
D.1 Detailed Ablation Results
To analyze the impact of each component, we systematically deactivate individual modules and observe the resulting performance (Table 10). In summary, deactivating the Adaptive Relabeling module leads to a significant decrease in known class detection (mAP) while maintaining a stable U-Recall. On the other hand, the SuperClass Regularizer and Hyperbolic Distance modules are responsible for the U-Recall performance. These modules play a crucial role in achieving a balance between known class detection (mAP) and unknown class retrieval (U-Recall) maintaining an equilibrium in the detection process.
Task 1 Task 2 Task 3 Task 4
U-Recall ($\uparrow$) mAP($\uparrow$) U-Recall ($\uparrow$) mAP ($\uparrow$) U-Recall ($\uparrow$) mAP ($\uparrow$) mAP ($\uparrow$)
Current Previous Current Both Previous Current Both Previous Current Both
Hyp-OW (Ours) 34.9 49.9 47.5 42.0 49.0 45.5 55.2 44.4 44.1 44.3 42.8 47.0 43.9
w/ Cosine Distance ($c=0$) 32.8 $\boldsymbol{(-2.1)}$ 49.0 $\boldsymbol{(-0.9)}$ 46.4 $\boldsymbol{(-1.1)}$ 42.0 48.8 45.4 55.4 43.6 42.3 43.2(1.$\boldsymbol{(-1.1)}$ 42.4 45.2 43.1
w/o SuperClass Regularizer ($\beta=0$) 32.0 $\boldsymbol{(-2.9)}$ 50.0 47.1 41.3 48.9 45.1 52.9 $\boldsymbol{(-2.3)}$ 43.8 43.5 43.7 42.8 45.9 43.5
w/o Adaptive Relabeling 34.7 41.2 $\boldsymbol{(-8.7)}$ 47.6 36.2 41.5 38.9 $\boldsymbol{(-6.6)}$ 54.1 $\boldsymbol{(-1.1)}$ 36.5 36.6 36.5 $\boldsymbol{(-7.8)}$ 34.5 37.9 36.1 $\boldsymbol{(-7.8)}$
Table 10: Impact of each component of Hyp-OW on Hierachichal Split.
Deactivating Relabeling Module:
To asses the impact of this module, we adopt PROB’s (Zohar et al., 2022) methodology where every unmatched $\mathbf{q^{\textit{u}}}$ is labelled as unknowns.
Deactivating SuperClass Regularizer:
This is done by setting $\beta=0$
Deactivating Hyperbolic Embedding:
This is done by setting $c=0$
D.2 Effectiveness of our Semantic Adaptive Relabeling Scheme
We now provide insights into the effectiveness of our Semantic Distance-based Relabeling Scheme. Figure 10 and  11 show the heatmap distance between known (left) and unknown items (bottom) for Task 1 in the OWOD and Hierarchical Split, respectively. Lighter colors indicate lower distances and higher similarity between known and unknown classes.
We can observe a significant similarity between known and unknown classes belonging to the same category, as highlighted by teal hatched boxes in both heatmaps. For example, in the OWOD Split, the animal classes from the knowns exhibit high similarity with the unknown classes. This is observed by comparing the classes on the left side of the heatmap, such as bird, cat, and cow, with the classes at the bottom, such as elephant and bear. In the Hierarchical Split, classes from the food category also demonstrate high similarity (see banana, apple versus cake and orange).
These findings emphasize the effectiveness of our Hyperbolic Distance-based Relabeling Scheme in capturing and leveraging the hierarchical and semantic relationships between known and unknown classes.
Figure 10: Hyperbolic Knowns-Unknowns Distance Heatmap: In the heatmap, lighter colors indicate higher similarity. We observe that classes belonging to the same category, such as ’car’ and ’truck’ or ’bed’ and ’sofa’, exhibit higher similarity.
Figure 11: Hyperbolic Knowns-Unknowns Distance Heatmap: Lighter colors indicate higher similarity. We can observe that classes from the same category, such as ’skis’ and ’snowboard’ or ’toaster’ and ’microwave’, exhibit higher similarity.
D.3 SuperClass Regularizer
Figure 12 illustrates the Hyperbolic Distance between each SuperClass (left) and class (bottom) with (top) and without (bottom) our Regularizer. The inclusion of the Regularizer results in a wider range of values (from $0.7$ to $2.30$) compared to the absence of the Regularizer, where the range is narrower (from $0.78$ to $1.2$). This demonstrates that our Regularizer effectively pushes classes from different categories apart while bringing classes from the same category closer together. Additionally, without the Regularizer, the values are compressed, resulting in a squeezed inter-category distance, as evidenced by the lighter colors and weaker color contrast in the top plot. T-sne visualization ( Figure 13) shows a qualitative visualization of this phenomenon.
Figure 12: Hyperbolic SuperClass-Class Heatmap Distance.
Figure 13: t-SNE plot of the learned class representations, with colors representing their respective categories.
D.4 Cosine versus Hyperbolic Distance
Figure 14 displays a t-SNE plot of the learned representation, where each class is color-coded based on its category. While both representations may initially appear scattered, the hyperbolic (right) distance-based representation exhibits improved grouping of classes from the same category. On the other hand, the cosine distance-based representation (left) lacks clear clustering properties between categories, as seen in the appliance category (black), animal category (light green), and furniture category (purple).
These observations highlight the advantage of utilizing hyperbolic distance for capturing meaningful category hierarchy relationship in the embedding space.
Figure 14: t-SNE plot of the learned class embedding where each color represents the class’ category.
D.5 Impact of Curvature Coefficient c
To investigate the influence of the curvature coefficient on the hyperbolicity of the embedding space, we conducted experiments using different values of $c$. Our focus was on Task 1 of the Hierarchical Split, where we assessed the model’s performance for four values of $c$: $0.0$, $0.1$, $0.2$, and $0.5$. The results are presented in Table 11. Notably, the curvature coefficient primarily impacts the U-Recall metric, while the mAP metric remains relatively stable. This observation is expected since the curvature coefficient $c$ directly influences the learned hierarchical structure (hence the Semantic Similarity Distance), which is the cornerstone of our Adaptive Relabeling Scheme for unknown retrieval. Among the tested values, we found that $c=0.1$ yielded the optimal performance, consistent with previous findings in the literature (Ermolov et al., 2022; Khrulkov et al., 2020b; Yue et al., 2023).
Hierarchical Split - Task 1
U-Recall ($\uparrow$) mAP($\uparrow$)
$c=0$ (Cosine Distance) 32.8 49.0
$c=0.1$ (Hyp-OW ) 34.9 49.9
$c=0.2$ 33.3 49.5
$c=0.5$ 32.3 49.8
Table 11: Impact of curvature coefficient $c$.